If the area of a sector of a circle is 8p and with central angle 45º, find the radius of the circle.


1
Expert's answer
2021-01-20T03:12:43-0500

"A=\pi r^2*\frac{45}{360}"

"\pi r^2*\frac{45}{360}=8\pi"

"r^2=\frac{8*360}{45}=64"

"r=\sqrt{64}=8"